(b) Find the asymptotes of the hyperbola
x
2 − y
2 + 2x + y + 9 = 0.
1
Expert's answer
2020-07-09T20:10:29-0400

x2-y2+2x+y+9=0

x2+2x-y2+y= -9

-(x2+2x)/9 + (y2-y)/9 =1

-{(x+1)2 -1}/9 + {y - 1/2)2 -1/4}/9 = 1+1/9+1/36

{(y - 1/2)2}/9 - ((x+1)2}/9 =1

now we got the value of thye centre of this hyperbola (h,k) = (-1 , 1/2)

now , the equation of the two asymptote of this hyperbola are

(1) y = 1/2 + x-(-1)

y = 1/2 +1 +x

y = x + 3/2

(2) y = 1/2 - x+(-1)

y = 1/2 -x -1

y = -x - 1/2